# Wireless Communication for Low-Altitude Economy with UAV Swarm Enabled Two-Level Movable Antenna System

### v1

@@ IV-B1 Optimization of ${\bf V}$ With Given ${\bf Q}$ and ${\bf B}$
$\mathop{\max}\limits_{\_}{{b_{\_}{l,m}},\rho,\left\{{{\eta_{\_}k},{\mu_{\_}k}}\right\}_{\_}{k=1}^{K}}\ \ \rho$ (51)
${\rm{s.t.}}$ $\ {{{\eta_{\_}k}-{\mu_{\_}k}}}\geq\rho\ln 2,\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{\_}{i=1}^{K}{{P_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}\geq{\sigma^{2}}{e^{{\eta_{\_}k}}},\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{\_}{i=1,i\neq k}^{K}{{P_{\_}i}{{\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|}^{2}}}\leq{\sigma^{2}}{e^{{\mu_{\_}k}}},\ \forall k\in{\cal K},$
${b_{\_}{l,m}}\in{\tilde{\cal C}},$
${\left|{{b_{\_}{l,m}}-{b_{\_}{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\_}{\min}^{2},\ \forall m^{\prime}\neq m.$
Let ${{\bf{\tilde{h}}}_{\_}{i,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ and ${{\bf{\tilde{v}}}_{\_}{i,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ denote the resulting channel and beamforming vector after removing ${h_{\_}{i,l,m}}$ and ${v_{\_}{i,l,m}}$, respectively, where ${h_{\_}{i,l,m}}$ and ${v_{\_}{i,l,m}}$ are the $\left({\left({l-1}\right)M+m}\right)$-th elements of ${{{\bf{h}}_{\_}i}}$ and ${{{\bf{v}}_{\_}k}}$, respectively. The expression ${f_{\_}{k,i}}={\left|{{\bf{v}}_{\_}k^{H}{{\bf{h}}_{\_}i}}\right|^{2}}$ can be expressed as
${f_{\_}{k,i}}=\underbrace{2\left|{{c_{\_}{k,i,l,m}}}\right|{\mathop{\rm Re}\nolimits}\left\{{{e^{{\rm{j}}\left({\frac{{2\pi}}{\lambda}{\Phi_{\_}i}{b_{\_}{l,m}}+\angle{c_{\_}{k,i,l,m}}}\right)}}}\right\}}_{\_}{{g_{\_}{k,i,l,m}}}+$ (52)
$\underbrace{{{\left|{{\alpha_{\_}i}}\right|}^{2}}{{\left|{{v_{\_}{k,l,m}}}\right|}^{2}}+{\bf{\tilde{h}}}_{\_}{i,l,m}^{H}{{{\bf{\tilde{v}}}}_{\_}{k,l,m}}{\bf{\tilde{v}}}_{\_}{k,l,m}^{H}{{{\bf{\tilde{h}}}}_{\_}{i,l,m}}}_{\_}{{{\bar{g}}_{\_}{k,i,l,m}}},$
where ${c_{\_}{k,i,l,m}}\triangleq{\bf{\tilde{h}}}_{\_}{i,l,m}^{H}{{{\bf{\tilde{v}}}}_{\_}{k,l,m}}v_{\_}{k,l,m}^{*}{\alpha_{\_}i}\times$ ${e^{{\rm{j}}\frac{{2\pi}}{\lambda}{\bm{\kappa}}_{\_}i^{T}\left({{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}r}}\right)}}$, and ${{{\bar{g}}_{\_}{k,i,l,m}}}$ is independent of ${{b_{\_}{l,m}}}$. Thus, problem (51) can be solved similar to (30).
The main procedures for solving problem (P1) are summarized in Algorithm 2. The complexity for obtaining the receive beamforming in step 3 is ${\cal O}({K{(LM)^{3}}})$. From step 4 to step 6, the complexity is approximately ${\cal O}({I_{\_}1}L{{({2K})}^{3}})$, where $I_{\_}1$ denotes the maximum number of iterations required by SCA for convergence. The complexity from step 7 to step 11 is approximately given by ${\cal O}({I_{\_}2}L(M-1){{({2K})}^{3}})$, where $I_{\_}2$ denotes the maximum number of iterations to converge required by step 9. As a result, the total computational complexity of Algorithm 2 is ${\cal O}({{I_{\_}3}K{(LM)^{3}}+{I_{\_}2}{I_{\_}3}L(M-1){({2K})}^{3}})$, where $I_{\_}3$ denotes the number of iterations required by the alternating optimization for convergence.
@@
1:  Initialize ${{\bf{Q}}^{\left(j\right)}}$, ${{\bf{B}}^{\left(j\right)}}$, and $\{{{\bf{v}}_{\_}k^{\left(j\right)}}\}$ randomly, and let $j=0$.
@@
3:     For given ${{\bf{Q}}^{\left(j\right)}}$ and ${{\bf{B}}^{\left(j\right)}}$, obtain the optimal beamforming $\{{{\bf{v}}_{\_}k^{\left({j+1}\right)}}\}$.
@@
5:      Obtain $\{{\bf{q}}_{\_}l^{\left({j+1}\right)}\}$ similar to problem (39), given  $\{{\bf{q}}_{\_}1^{\left({j+1}\right)},\cdots,{\bf{q}}_{\_}{l-1}^{\left({j+1}\right)},{\bf{q}}_{\_}{l+1}^{\left(j\right)},\cdots,{\bf{q}}_{\_}L^{\left(j\right)}\}$, ${{\bf{B}}^{\left(j\right)}}$,  and $\{{{\bf{v}}_{\_}k^{\left({j+1}\right)}}\}$.
@@
9:       Obtain ${b_{\_}{l,m}^{\left({j+1}\right)}}$ given ${{\bf{Q}}^{\left({j+1}\right)}}$, $\{b_{\_}{1,2}^{\left({j+1}\right)},\cdots,$  $b_{\_}{l,m-1}^{\left({j+1}\right)},b_{\_}{l,m+1}^{\left(j\right)},\cdots b_{\_}{L,M}^{\left(j\right)}\}$, and $\{{{\bf{v}}_{\_}k^{\left({j+1}\right)}}\}$.
@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{\_}0}=1$ m is ${\beta_{\_}0}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\_}{\min}}=1$ m. The movable region of UAVs is specified by ${x_{\_}l}\in\left[{-30,30}\right]$ m, ${y_{\_}l}\in\left[{-30,30}\right]$ m, and ${z_{\_}l}\in\left[{85,115}\right]$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\_}{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and $200$ m, respectively. The transmit power of each UE is $P_{\_}k=10$ dBm.
@@ V-A Single-Antenna UAV
Fig. 4 shows the minimum achievable rate versus the transmit power of each UE for the two-UE communication, by considering the UE directions $\left({{\theta_{\_}1},{\phi_{\_}1}}\right)=\left({{{60}^{\circ}},{{30}^{\circ}}}\right)$ and $\left({{\theta_{\_}2},{\phi_{\_}2}}\right)=\left({{{30}^{\circ}},{{0}^{\circ}}}\right)$, respectively. For comparison, the benchmark scheme of circular array geometry is considered, i.e., the UAV swarm cooperatively forms a circular array, with their placement positions given by
${{\bf{q}}_{\_}l}{\rm=}{{\bf{q}}_{\_}r}+\left[R\cos\left({\frac{{2\pi}}{L}\left({l-1}\right)}\right),{R\sin\left({\frac{{2\pi}}{L}\left({l-1}\right)}\right),0}\right]^{T},\ \forall l,$ (53)
where $R$ denotes the radius of circular array geometry and is set as $R=7.5$ m. With the circular array geometry, the MMSE, ZF and MRC beamforming schemes are respectively considered. It is observed that the proposed UAV swarm enabled MA system achieves an IUI-free communication, thanks to the flexible placement position optimization. Besides, the proposed scheme significantly outperforms the benchmark schemes of circular array geometry with MMSE, ZF, and MRC beamforming. This is expected since the proposed scheme is able to completely orthogonalize the channels of two UEs, while achieving the full beamforming gain at each UE. Specifically, Fig. 5 shows the SNR loss factor versus the transmit power for UE 1. It is observed that the proposed scheme always enjoys a zero SNR loss factor. By contrast, the circular array geometry with MMSE and MRC beamforming schemes experience increased SNR loss factors as the transmit power increases, as a result of suffering from more severe IUI, especially for the MRC beamforming scheme.
@@
Fig. 7 shows the minimum achievable rate versus the number of UEs, $K$. It is observed that the performance of both schemes decrease as the number of UEs increases. This is because more UEs will cause a severer IUI issue, especially when the number of UAVs/antennas is smaller than that of UEs. It is also observed that the minimum achievable rate of the proposed UAV swarm enabled MA system always surpasses the benchmark scheme of circular array geometry, thanks to the flexible UAV placement position adjustment to significantly reduce the channel correlation and thus the IUI among UEs.
@@ V-B Multi-Antenna UAV
Last, Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the UE directions are the same as for Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (20) and (21); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (53); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (20) and (21) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is bounded by the IUI-free communication from above.
@@ Appendix A Proof of Theorem 1
By substituting ${{\bf{q}}_{\_}l}={{\bf{q}}_{\_}1}+\left({l-1}\right)\times$ $\frac{{\left({{\varsigma}+1/L}\right)\lambda}}{{{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}^{2}}}}\left({{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right)$ into (19), where $\varsigma$ is a common coefficient to all the UAVs, given by
@@
$\left\|{{{\bf{q}}_{\_}l}-{{\bf{q}}_{\_}{l^{\prime}}}}\right\|=\left\|{\left({l-l^{\prime}}\right)\left({\varsigma+1/L}\right)\lambda\frac{{\left({{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right)}}{{{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}^{2}}}}}\right\|$ (55)
$\geq\left\|{\left({\varsigma+1/L}\right)\lambda\frac{{\left({{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right)}}{{{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}^{2}}}}}\right\|=\frac{{\lambda\left|{\varsigma+1/L}\right|}}{{\left\|{{{\bm{\kappa}}_{\_}k}-{{\bm{\kappa}}_{\_}{k^{\prime}}}}\right\|}}.$
To satisfy the collision avoidance constraint for UAVs, a feasible solution of $\varsigma$ is thus given by (21). This completes the proof of Theorem 1.

### v2

@@ IV-B1 Optimization of $\bf V$} With Given $\bf Q$} and $\bf B$}
$\mathop{\max}\limits_{{b_{l,m}},\rho,\left\{{{\eta_{k}},{\mu_{k}}}\right\}_{k=1}^{K}}\ \ \rho$ (50)
${\rm{s.t.}}$ $\ {{{\eta_{k}}-{\mu_{k}}}}\geq\rho\ln 2,\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{i=1}^{K}{{P_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}\geq{\sigma^{2}}{e^{{\eta_{k}}}},\ \forall k\in{\cal K},$
${\sigma^{2}}+\sum\limits_{i=1,i\neq k}^{K}{{P_{i}}{{\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|}^{2}}}\leq{\sigma^{2}}{e^{{\mu_{k}}}},\ \forall k\in{\cal K},$
${b_{l,m}}\in{\tilde{\cal C}}$,
${\left|{{b_{l,m}}-{b_{l,m^{\prime}}}}\right|^{2}}\geq\tilde{d}_{\min}^{2},\ \forall m^{\prime}\neq m.$
Let ${{\bf{\tilde{h}}}_{i,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ and ${{\bf{\tilde{v}}}_{k,l,m}}\in{{\mathbb{C}}^{\left({LM-1}\right)\times 1}}$ denote the resulting channel and beamforming vector after removing $h_{i,l,m}$} and $v_{i,l,m}$}, respectively, where $h_{i,l,m}$} and $v_{i,l,m}$} are the $\left({\left({l-1}\right)M+m}\right)$-th elements of ${\bf{h}}_{i}$}} and ${\bf{v}}_{k}$}}, respectively. The term ${f_{k,i}}={\left|{{\bf{v}}_{k}^{H}{{\bf{h}}_{i}}}\right|^{2}}$ can be further expressed as
${f_{k,i}}=\underbrace{2\left|{{c_{k,i,l,m}}}\right|{\mathop{\rm Re}\nolimits}\left\{{{e^{-{\rm{j}}\left({\frac{{2\pi}}{\lambda}{\Phi_{i,l}}{b_{l,m}}-\angle{c_{k,i,l,m}}}\right)}}}\right\}}_{{g_{k,i,l,m}}}+$ (51)
$\underbrace{{{\left|{{\alpha_{i}}}\right|}^{2}}{{\left|{{v_{k,l,m}}}\right|}^{2}}+{\bf{\tilde{h}}}_{i,l,m}^{H}{{{\bf{\tilde{v}}}}_{k,l,m}}{\bf{\tilde{v}}}_{k,l,m}^{H}{{{\bf{\tilde{h}}}}_{i,l,m}}}_{{{\bar{g}}_{k,i,l,m}}},$
where ${c_{k,i,l,m}}\triangleq{\bf{\tilde{h}}}_{i,l,m}^{H}{{{\bf{\tilde{v}}}}_{k,l,m}}v_{k,l,m}^{*}{\alpha_{i}}{e^{-{\rm{j}}\frac{{2\pi}}{\lambda}{r_{i,l}}}}$, and ${\bar{g}}_{k,i,l,m}$}} is independent of $b_{l,m}$}}. Thus, problem (50) can be solved similar to (29).
The main procedures for solving problem (P1) are summarized in Algorithm 2. The complexity for obtaining the receive beamforming in step 3 is ${\cal O}({K{(LM)^{3}}})$. From step 4 to step 6, the complexity is approximately ${\cal O}({I_{1}}L{{({2K})}^{3}})$, where $I_{1}$ denotes the maximum number of iterations required by SCA for convergence. The complexity from step 7 to step 11 is approximately given by ${\cal O}({I_{2}}L(M-1){{({2K})}^{3}})$, where $I_{2}$ denotes the maximum number of iterations to converge required by step 9. As a result, the total computational complexity of Algorithm 2 is ${\cal O}({{I_{3}}K{(LM)^{3}}+{I_{2}}{I_{3}}L(M-1){({2K})}^{3}})$, where $I_{3}$ denotes the number of iterations required by the alternating optimization for convergence.
@@
1: Initialize ${\bf{Q}}^{\left(0\right)}$}, ${\bf{B}}^{\left(0\right)}$}, and $\{{{\bf{v}}_{k}^{\left(0\right)}}\}$ randomly, and let $j=0$.
@@
3:  For given ${\bf{Q}}^{\left(j\right)}$} and ${\bf{B}}^{\left(j\right)}$}, obtain the optimal beamforming $\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\}$.
@@
5:   Obtain $\{{\bf{q}}_{l}^{\left({j+1}\right)}\}$ similar to problem (38), given  $\{{\bf{q}}_{1}^{\left({j+1}\right)},\cdots,{\bf{q}}_{l-1}^{\left({j+1}\right)},{\bf{q}}_{l}^{\left(j\right)},\cdots,{\bf{q}}_{L}^{\left(j\right)}\}$, ${\bf{B}}^{\left(j\right)}$},  and $\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\}$.
@@
9:    Obtain $b_{l,m}^{\left({j+1}\right)}$} given ${\bf{Q}}^{\left({j+1}\right)}$}, $\{b_{1,2}^{\left({j+1}\right)},\cdots,$  $b_{l,m-1}^{\left({j+1}\right)},b_{l,m}^{\left(j\right)},\cdots b_{L,M}^{\left(j\right)}\}$, and $\{{{\bf{v}}_{k}^{\left({j+1}\right)}}\}$.
@@ V Numerical Results
In this section, numerical results are provided to verify the performance of the proposed low-altitude UAV swarm enabled MA system. The channel power at the reference distance of ${d_{0}}=1$ m is ${\beta_{0}}=-61.4$ dB, and the noise power is ${\sigma^{2}}=-94$ dBm. The minimum distance to avoid the collision among UAVs is ${d_{\min}}=1$ m. Moreover, the minimum distance to avoid mutual coupling between adjacent MA elements is ${{\tilde{d}}_{\min}}=\lambda/2$. The (local) movable region of MAs at each UAV is $\left[{0,D}\right]$, with $D=20\lambda$. Unless otherwise stated, $K=3$ UEs are uniformly distributed in a circular area with the center and radius being $\left[{0,0,0}\right]^{T}$ m and $600$ m, respectively. The transmit power of each UE is $P_{k}=20$ dBm. Without otherwise stated, the movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-50$ m, ${x_{\max}}={y_{\max}}=50$ m, ${z_{\max}}=40$ m, and $H=500$ m.
@@ V-A Single-Antenna UAV
Fig. 4 shows the minimum achievable rate versus the transmit power of each UE for the two-UE communication, by considering the UE directions $\left({{\theta_{1}},{\phi_{1}}}\right)=\left({{{30}^{\circ}},{{60}^{\circ}}}\right)$ and $\left({{\theta_{2}},{\phi_{2}}}\right)=\left({{{0}^{\circ}},{{0}^{\circ}}}\right)$, respectively. The movable region of UAVs is specified by ${{\bar{x}}_{l}}\in\left[{{x_{\min}},{x_{\max}}}\right]$ m, ${{\bar{y}}_{l}}\in\left[{{y_{\min}},{y_{\max}}}\right]$ m, and ${{\bar{z}}_{l}}\in\left[{H,H+z_{\max}}\right]$ m, with ${x_{\min}}={y_{\min}}=-20$ m, ${x_{\max}}={y_{\max}}=20$ m, ${z_{\max}}=10$ m, and $H=1000$ m. For comparison, the benchmark scheme of circular array geometry is considered, i.e., the UAV swarm cooperatively forms a circular array, with their placement positions given by
${{\bf{\bar{q}}}_{l}}{\rm=}{{\bf{q}}_{r}}+\left[R\cos\left({\frac{{2\pi}}{L}\left({l-1}\right)}\right),{R\sin\left({\frac{{2\pi}}{L}\left({l-1}\right)}\right),0}\right]^{T},\ \forall l,$ (52)
where $R$ denotes the radius of circular array geometry and is set as $R=\left({{x_{\max}}-{x_{\min}}}\right)/2=\left({{y_{\max}}-{y_{\min}}}\right)/2$. With the circular array geometry, the MMSE, ZF and MRC beamforming schemes are respectively considered. It is observed that the proposed UAV swarm enabled MA system achieves an IUI-free communication, thanks to the flexible placement position optimization. Besides, the proposed scheme outperforms the benchmark schemes of circular array geometry with MMSE, ZF, and MRC beamforming. This is expected since the proposed scheme is able to completely orthogonalize the channels of two UEs, while achieving the full beamforming gain at each UE. Specifically, Fig. 5 shows the SNR loss factor versus the transmit power for UE 1. It is observed that the proposed scheme always enjoys a zero SNR loss factor. By contrast, the circular array geometry with MMSE and MRC beamforming schemes experience increased SNR loss factors as the transmit power increases, as a result of suffering from more severe IUI, especially for the MRC beamforming scheme.
@@
Fig. 7 shows the minimum achievable rate versus the number of UEs, $K$. It is observed that the performance of both schemes decrease as the number of UEs increases. This is because more UEs will cause a severer IUI issue, especially when the number of UAVs/antennas is smaller than that of UEs, e.g., $K>4$. It is also observed that the minimum achievable rate of the proposed UAV swarm enabled MA system always surpasses the benchmark scheme of circular array geometry, thanks to the flexible UAV placement position adjustment to significantly reduce the channel correlation and thus the IUI among UEs.
@@ V-B Multi-Antenna UAV
Last, Fig. 9 studies the impact of the normalized movable region size $D/\lambda$ on the minimum achievable rate. For $K=2$, the transmit power of each UE is ${P_{k}}=35$ dBm, the UE directions are the same as for Fig. 4, and the following three schemes are considered: 1) Placement optimization with USA: each UAV is equipped with a USA and UAV swarm adopts the placement given in (21) and (22); 2) Circular array geometry with USA: each UAV is equipped with a USA and UAV swarm adopts the circular geometry given in (52); 3) Circular array geometry with FPA: Similar to 2), but using FPA instead USA. For USA, the sparsity level can be dynamically adjusted. It is firstly observed from Fig. 9(a) that the scheme of placement optimization with USA directly achieves the IUI-free communication. This is because the placement given in (21) and (22) orthogonalizes the channels of two UEs. On the other hand, as the normalized movable region size increases, the scheme of circular array geometry with USA also achieves the IUI-free communication, thanks to the extra DoF of sparsity level adjustment for completely eliminating the IUI. Moreover, for $K=3$, it is observed from Fig. 9(b) that the performance of two FPA schemes remain unchanged as the normalized movable region size increases, as expected. By contrast, the performance of the proposed UAV swarm enabled MA system first increases and ultimately converges as the normalized movable region size increases. This is mainly attributed to the fact that with an enlarged movable region size, MA array has a larger spatial DoF to create favorable channels for performance improvement. However, this does not mean that the performance gain of UAV swarm enabled MA system over other benchmark schemes will continuously increase, since it is bounded by the IUI-free communication from above.
@@ VI Conclusion
In addition, there are many important research directions that could be pursued in the future. For instance, the channel modeling incorporating the array misalignment, orientation jitter, and synchronization issues, as well as the extension to multi-path channels, are worthwhile to investigate. Besides achievable rate, energy consumption and latency are the important performance metrics for UAV swarm enabled MA system, while designing the energy-efficient and low-latency UAV system needs more in-depth studies. Moreover, the flight control mechanisms of UAV swarm to guarantee the safety and mission execution are important to investigate in the future. Last, how to integrate IRS into UAV swarm enabled MA systems is an interesting direction that deserves further investigation.
@@ Appendix A Proof of Theorem 1
By substituting ${{\bf{q}}_{l}}={{\bf{q}}_{1}}+\left({l-1}\right)\times$ $\frac{{\left({{\varsigma}+1/L}\right)\lambda}}{{{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}^{2}}}}\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)$ into (20), where $\varsigma$ is a common coefficient to all the UAVs, given by
@@
$\left\|{{{\bf{q}}_{l}}-{{\bf{q}}_{l^{\prime}}}}\right\|=\left\|{\left({l-l^{\prime}}\right)\left({\varsigma+1/L}\right)\lambda\frac{{\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)}}{{{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}^{2}}}}}\right\|$ (54)
$\geq\left\|{\left({\varsigma+1/L}\right)\lambda\frac{{\left({{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right)}}{{{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}^{2}}}}}\right\|=\frac{{\lambda\left|{\varsigma+1/L}\right|}}{{\left\|{{{\bm{\kappa}}_{k}}-{{\bm{\kappa}}_{k^{\prime}}}}\right\|}}.$
To satisfy the minimum safe distance constraint for UAVs, a feasible solution of $\varsigma$ is thus given by (22). This completes the proof of Theorem 1.